The sum of the ages of Petra and her mother is 48. Her mother is 12 years more than twice as old as Petra. How old are Petra and her mother?
Answer by feliz1965(151)

Petra's age = 2x
Petra mom's age = 2x + 12
Their sum is 48.
2x + 2x + 12 = 48
4x + 12 = 48
4x = 48 - 12
4x = 36
x = 36/4
x = 9
Petra is 2(9) = 18 years old.
Petra's mom is 18 + 12 = 30 years old.